OpenStudy (anonymous):

What is the solution to the following equation? 6(x - 3) = 2(13 - x)

OpenStudy (anonymous):

6x-18 = 26-2x, bring the x's to one side and the other numbers to the other side so the equation will look like : 6x+2x =26+18 8x = 44 divide both sides by 8 x = 44/8 then simplify x = 11/4